Phyllis invested 38000 dollars, a portion earning a simple interest rate of 4 percent per year and the rest earning a rate of 6 percent per year. After one year the total interest earned on these investments was 1860 dollars. How much money did she invest at each rate?
Please help me, I'm having a tough time coming up with an equation and finding the answer! Thanks so much for you time.

Answer by richwmiller(17219)

Total amount of money invested: $38000
x+y=38000,
Total yearly interest for the two accounts is: $1860
0.04*x+0.06*y=1860
x=38000-y
Substitute for x
0.04*(38000-y)+0.06*y=1860
Multiply out
1520-0.04y+0.06*y=1860
Combine like terms.
0.02*y=340
Isolate y
y=$17000.00 at 6%
x=38000-y
Calculate x
x=$21000.00 at 4%
Check
0.04*21000+0.06*17000=1860
840+1020=1860
1860=1860
If this statement is TRUE and neither amount is negative then all is well